electric field on infinite line charge due to finite line charge of length L and distance between charges is a.

Grade:12

2 Answers

Eshan
5 years ago
From Newton’s third law of motion, or since the center of mass of the body does not accelerate(absence of external force), the force acting on both the bodies is equal in magnitude. Hence the force acting on inifinite line charge due to finite line charge is equal to that by infinite line charge on finite line charge=qE=
(\lambda_1L)(\dfrac{\lambda_2}{2\pi\epsilon_0 a})=\dfrac{\lambda_1\lambda_2L}{2\pi\epsilon_0 a}
Eshan
5 years ago
From Newton’s third law of motion, or since the center of mass of the body does not accelerate(absence of external force), the force acting on both the bodies is equal in magnitude. Hence the force acting on inifinite line charge due to finite line charge is equal to that by infinite line charge on finite line charge=qE=
(\lambda_1L)(\dfrac{\lambda_2}{2\pi\epsilon_0 a})=\dfrac{\lambda_1\lambda_2L}{2\pi\epsilon_0 a}